# Standardizing the Measurement of Text Diversity: A Tool and Comparative Analysis

### v2

@@ Abstract
The diversity across outputs generated by LLMs shapes perception of their quality and utility. High lexical diversity is often desirable, but there is no standard method to measure this property. Templated answer structures and “canned” responses across different documents are readily noticeable, but difficult to visualize across large corpora. This work aims to standardize measurement of text diversity. Specifically, we empirically investigate the convergent validity of existing scores across English texts, and we release diversity, an open-source Python package [^1] https://pypi.org/project/diversity/ for measuring and extracting repetition in text. We also build a platform [^2] https://ai-templates.app based on diversity for users to interactively explore repetition in text. We find that fast compression algorithms capture information similar to what is measured by slow-to-compute $n$-gram overlap homogeneity scores. Further, a combination of measures—compression ratios, self-repetition of long $n$-grams, and Self-BLEU—are sufficient to report, as they have low mutual correlation with each other.
@@ A Tool and Comparative Analysis
Chantal Shaib1thanks: Partial work completed while at Adobe Research.  Joe Barrow3footnotemark:   Jiuding Sun1  Alexa F. Siu2 Byron C. Wallace1  Ani Nenkova2 1Northeastern University, 2Adobe Research, 3Pattern Data shaib.c@northeastern.edu
@@ 1 Introduction
Figure 1: Part-of-speech pattern visualization with the diversity app. Users upload a corpus, and the UI returns frequent part-of-speech sequences. Users can navigate through different pattern lengths of $n=(2,10)$.
@@
Evaluation of LLM-generated texts is typically done with respect to accuracy or factuality, e.g., as measured via entailment (Tang et al., 2023), or text quality aspects such as coherence and fluency (e.g., estimated using LLMs as evaluators as in Liu et al. 2023). For tasks where reference summaries are available, the similarity of generated outputs to these is also often measured (e.g., using ROUGE scores; Lin and Och, 2004). A complementary dimension of model performance is diversity, or how much “boilerplate” content is repeated across LLM outputs. For example, prompted to summarize news articles, Llama-2 (Touvron et al., 2023b) tends to generate text that begins with “The article discusses that…", “The article notes that…". Table 1 illustrates example repetitions from various models over a news summarization dataset.
Token Repetition Text
"The article also notes that…" (41/500 )
"The article also mentions that…" (17/500)
"The article notes that the…" (15/500)
"The article discusses the…" (15/500)
"The article also mentions that…" (28/500)
"The article is about the…" (19/500)
"According to the article, a…" (16/500)
Pattern-Matched Text
DT NN VBZ DT JJ NN (84/500)
"The article discusses the recent debate…"
"The book provides a helpful guide…"
"The article discusses the controversial penalty…"
"The article discusses the illicit market…"
NNP NNP DT JJ NN VBD (37/500)
"Christopher Barry, a 53-year-old man, was…"
"Charles Collins, a 28-year-old man, saved…"
"Damian Parks, a 22-year-old student, went…"
"Lynn Fast, a 21-year-old mother, claimed…"
Table 1: Exact text match and repeated part-of-speech patterns across 500 LLM-generated summaries of news articles from CNN/DM. The number of times the pattern occurs is reported parenthetically.
There is a distinct lack of standardization in reporting diversity in machine learning datasets (Zhao et al., 2024). Evaluating diversity is complicated by a lack of tools enabling quantification and exploration of repetitive text in large corpora. This work primarily aims to standardize such reporting, focusing on text-based corpora. We introduce diversity, an open-source Python package for evaluating diversity.1 In addition to this package, we provide a standalone web-based tool that allows users to visualize repetition in their corpus.2 This UI provides an intuitive, efficient tool for text analysis, allowing users to: (i) View repetitive text and part-of-speech $n$-grams of varying lengths; (ii) Quickly compute diversity metrics, and; (iii) Interactively highlight and match repetition in documents. To facilitate future exploration and measurement of lexical diversity, both package333 https://github.com/cshaib/diversity and UI code [^4] https://github.com/cshaib/diversity_demo are available under the Apache 2.0 license.
To validate existing diversity metrics, we analyze them over English language outputs from several LLMs. This identifies a few practical, (mostly) independent scores that characterize repetition. We emphasize text length as an important confounder when assessing diversity, and examine diversity in downstream datasets such as instruction tuning. When measuring lexical diversity, many comparisons remain meaningful when accompanied with information about length, but absent this variable no reliable conclusions can be drawn.
In addition to assessing the information overlap of existing metrics, we show that compression ratio—a measure of document compression relative to original size—is a fast, easy to compute score that is sufficient to capture the information in all token/type ratio related alternatives. However, we also find that compression ratios (and all scores considered) are moderately to strongly correlated with text length, complicating interpretation.
Our contributions are as follows. (1) We introduce diversity, a Python package implementing diversity metrics. (2) We host and release source code to a user interface to explore repetition and diversity. (3) We evaluate the convergent validity of existing lexical diversity metrics and highlight compression ratios as efficient measures of diversity. (4) We demonstrate use cases of our package in NLP research via published work.
@@ 2 Related Work
Lack of diversity in text may result from repetition of lengthy strings or owe to subtle distributional patterns (Holtzman et al., 2019; Meister et al., 2022, 2023a). We focus on scores that aim to capture overt repetition across outputs, and leave for future work similar analysis of semantic and structural diversity scores (Bär et al., 2012; Shaib et al., 2024).
Conditional generation tasks such as image captioning and dialog summarization have offered observations regarding the diversity of produced texts. Prior work on both of these tasks has documented that models tend to repeat the same text for different contexts. Li et al. (2016) find that four phrases account for about a third of all turns produced by a conversational agent, and Devlin et al. (2015) report that more than half of automatically generated captions are repeated verbatim for different images. Self-repetition (Salkar et al., 2022), i.e., exact repetition of the same $n$-gram ($n\geq 4$) across different summaries is a practical way of measuring repetition for tasks where generated text is lengthier. In such cases exact matches rarely occur but repetition is common, especially relative to the training data (Wang et al., 2023a).
Above we discussed a variety of metrics but it is unclear which of these to use when, and how to efficiently visualize the implications of lower diversity in tasks that comprise only text. Further, prior work has shown that human judgments of diversity are difficult to reliably collect. Humans tend to implicitly conflate quality of text with its diversity, and it can be difficult to separate content and lexical diversity in these assessments (Tevet and Berant, 2021). Motivated by the lack of visualization platforms and the need to provide better methods for human evaluations of repetition in large datasets, we build an interactive tool that allows users to browse highlighted instances of “lower diversity” text, as shown in Figure 1.
@@ 2.1 A Smorgasbord of Text Diversity Scores
Scores used to measure diversity across a corpus of texts derive from two core ideas: Computing average similarity between pairs of outputs produced by the same model for different inputs, and computing variants of token/type ratio. The former are adapted from common approaches to reference-based text generation using standard measures of pairwise similarity; the latter track the diversity of vocabulary measured as the ratio of unique words to total words produced, with outputs from a model concatenated into a single text. We first describe each score, and then present insights regarding their mutual redundancy. We also consider run-times, which are lengthy for some metrics and may render them impractical for analysis of a large datasets. All scores are defined for a set of generated texts $D$, each conditioned on its respective input.
Self-BLEU
The quality of text in machine translation, summarization, and image captioning is often reported in terms of overlap with a reference text. This idea can be adapted to measure diversity across different outputs by using one generated text as a “reference” and measuring the similarity of other outputs against this. Self-BLEU measures similarity between all text pairs in $D$ using BLEU (Zhu et al., 2018). BLEU could be replaced with other similarity scores, e.g., ROUGE-L or BERTScore. These variants are called homogenization scores and have recently been used to compare the diversity of texts produced under several conditions (Padmakumar and He, 2023a).
Homogenization Score (ROUGE-L)
All homogenization scores calculate an aggregate similarity across pairs of examples (Equation 1). Here the similarity score of choice is ROUGE-L Lin and Och 2004, which quantifies overlap in terms of longest common sub-sequences between all pairs of text in a corpus instead of the fixed $n$-gram size used in other ROUGE variants:
@@
Homogenization Score (BERTScore)
This homogenization score uses BERTScore to measure similarity between documents in Equation 1. Unlike the other scores, it does not count the repetition of specific tokens, but instead uses BERT embeddings to (ideally) capture “semantic” similarity beyond verbatim $n$-gram matches.
Self-repetition Score
Self-repetition measures the tendency of LMs to repeat long $n$-grams across different outputs (Salkar et al., 2022).
@@
Moving Average Token-Type Ratio
The token-type ratio for a text is the unique token count divided by the total count of tokens. Moving Average Token Type Ratios (MATTRs) measures the lexical dynamics across a text which is insensitive to text length. This captures the repetition of a given word in segments of text and does not explicitly account for longer repeated sequences (Covington and McFall, 2010).
$N$-Gram Diversity Score
NGD extends the idea of token-type ratio to longer $n$-grams (Padmakumar and He, 2023a; Meister et al., 2023b; Li et al., 2023). It is defined as a ratio of unique $n$-gram counts to all $n$-gram counts:
$\text{NGD}(D)=\sum\limits_{n=1}^{4}\frac{\text{\# unique }n\text{-grams in }D\oplus}{\text{\# }n\text{-grams in }D\oplus}$ (3)
@@
Hypergeometric Distribution D
The probability of text under a Hypergeometric Distribution D (HD-D) is an another measure of lexical diversity (McCarthy and Jarvis, 2010). [^5] For both HD-D and MATTR, we use the implementation provided in the lexical-diversity package (https://pypi.org/project/lexical-diversity/). HD-D does not capture repetition of sub-sequences.
@@ 2.2 Compression Ratios for Diversity
Compression Ratios
The diversity scores introduced so far are all a function of the number of repeated substrings across outputs. We use gZip to compress the concatenated text of all outputs generated by a model. The compression ratio is then the ratio between the size of the compressed file to that of the original file. High compression ratios imply more redundancy:
@@
Part-of-Speech Compression Ratio
To capture repeated syntactic patterns, we also compute compression ratios for part-of-speech (POS) tag sequences. We use the NLTK POS tagger  [^6] https://www.nltk.org/api/nltk.tag.html and the Penn Treebank set of 36 tags.
@@ 3.1 Design of the Diversity Package
The diversity package uses NLTK (Bird and Loper, 2004) and SpaCY Honnibal et al. (2020) to tag text with parts of speech and extract $n$-grams. Users can install the package via pip (assuming Python 3.10+). They can then calculate various diversity metrics over a list of texts as follows:
1from diversity import compression_ratio,
2 homogenization_score,
3 ngram_diversity_score,
4 get_pos
5
6text = [
7"I enjoy walking with my cute dog for the rest of the day, ...",
8"I enjoy walking with my cute dog. The only time I felt ...",
9"I enjoy walking with my cute dog, and I really enjoy running..."
10]
11
12# compression ratios
13cr = compression_ratio(text, ’gzip’)
14cr_pos = compression_ratio(get_pos(text)[1], ’gzip’)
15
16# homogenization scores
17hs_rougel = homogenization_score(text, ’rougel’)
18hs_bert = homogenization_score(text, ’bertscore’)
19self_bleu = homogenization_score(text, ’bleu’)
20
21# other
22self_rep = self_repetition_score(text)
23nds = ngram_diversity_score(text, n=4)
Users can also extract repeated text based on part of speech ngrams with the extract_patterns and match_patterns functions.
1n = 5
2top_n = 100
3extract_patterns(text, n, top_n)
First, the user extracts patterns using the extract_patterns function by specifying the n-gram length to search for, and the top_n most repeated n-grams to return. This returns a Python dictionary where the keys are the part-of-speech n-grams and the values are the raw text n-grams matching those patterns. The pattern matches are based on the frequency seen across the entire dataset, i.e., a part-of-speech pattern is only a pattern if it appears in more than 2 texts in the original input. Default values consider the top 100 part-of-speech patterns (sorted by frequency).
@@
which returns a list of tuples containing the pattern in the first position, and the matched substring in the second position.
Many diversity metrics require pairwise comparisons. With larger datasets, this can become infeasible to compute (see Appendix B). We implement a few methods to increase efficiency: We cache to store already computed pairs, and on the website we chunk incoming data and compute pattern searches over batches.
3.2 Design of the Web UI
The diversity Web UI functionality is built around the Python package. All front-end design uses Vue.js and Tailwind CSS.
The UI offers the same functionality as the package, however allows users to measure and explore repetition without requiring any Python knowledge. Figure 2 shows the main pages of the site: users can begin by (A) either uploading their own text file for analysis or selecting one of the demo datasets provided on the site. Then, the user is prompted to select one of three types of analyses: either (B) to explore part-of-speech patterns, (C) to explore verbatim repeated text, or (D) to measure various diversity metrics of the dataset. Datasets are processed upon upload, and nothing is stored on the backend server aside from the existing demo datasets.
(B) Templates
The templates tab allows users to explore extracted part-of-speech $n$-grams in their selected dataset. The left-most column displays pattern length of $n=[2,10]$. The user can then scroll through all of the templates, select some or all, and see the highlighted text in the middle panel corresponding to the template. The templates are assigned a colour when selected to indicate the corresponding matched text. The right-most column provides a reference for all the part-of-speech tags from SpaCY.  [^7] https://spacy.io/usage/linguistic-features The default pattern length is set to $n=4$. Other lengths will load when selected.
(C) Exact Match
The exact matches tab allows a user to explore exact text matches in their dataset. The top provides two sliders: the left slider allows the user to set a string length to search for ($n=[2,10]$), and the right the minimum number of documents in which the string must appear ($n=[2,10]$). The minimum document occurrence slider defaults to 2. Once selected, the user can scroll through to see the repeated text in bold, and the full document text in which the string appears, as well as the number of documents.
(D) Diversity Metrics
The diversity metrics tabs reports the recommended metrics from our evaluation: Compression Ratio, POS Compression, Self-BLEU, Self-Repetition, and Homogenization with BERTScore. We display these values alongisde a guide to the metrics on the right-hand side.

### v3

@@ Abstract
The diversity across outputs generated by LLMs shapes perception of their quality and utility. High lexical diversity is often desirable, but there is no standard method to measure this property. Templated answer structures and “canned” responses across different documents are readily noticeable, but difficult to visualize across large corpora. This work aims to standardize measurement of text diversity. Specifically, we empirically investigate the convergent validity of existing scores across English texts, and release diversity, an open-source Python package [^1] https://pypi.org/project/diversity/ for measuring and extracting repetition in text. We also build a platform [^2] https://ai-templates.app based on diversity for users to interactively explore repetition in text. We find that fast compression algorithms capture information similar to what is measured by slow-to-compute $n$-gram overlap homogeneity scores. Further, a combination of measures—compression ratios, self-repetition of long $n$-grams, and Self-BLEU—are sufficient to report, as they have low mutual correlation with each other.
@@ A Tool and Comparative Analysis
Chantal Shaib1thanks: Partial work completed while at Adobe Research.    Venkata S Govindarajan4    Joe Barrow3footnotemark:     Jiuding Sun1 Alexa F. Siu2  Byron C. Wallace1  Ani Nenkova2 1Northeastern University, 2Adobe Research, 3Pattern Data, 4Ithaca College shaib.c@northeastern.edu
@@ 1 Introduction
LLM-generated texts are typically evaluated with respect to accuracy or factuality, e.g., as measured via entailment (tang-etal-2023-understanding), or text quality aspects such as coherence and fluency (e.g., estimated using LLMs as evaluators Liu et al. 2023). When reference summaries are available, the similarity of generated outputs to these is also often measured (e.g., via ROUGE; lin-och-2004-automatic). A complementary dimension of model performance is diversity, or how much “boilerplate” content is repeated across LLM outputs.
There is a distinct lack of standardization in reporting diversity in ML datasets (Zhao et al., 2024). We address this by introducing diversity, an open-source Python package for evaluating text diversity,1 along with a web-based UI that allows users to visualize repetition in their corpus,2 providing an intuitive, efficient tool for text analysis that permits: (i) Viewing repetitive text and Part-of-Speech $n$-grams; (ii) Quickly computing diversity metrics, and; (iii) Interactively highlighting and matching repetition in documents. Both the package [^3] https://github.com/cshaib/diversity and UI code [^4] https://github.com/cshaib/diversity_demo are open-sourced under the Apache 2.0 license.
We run existing diversity metrics over English language outputs from several LLMs to identify a few (mostly) independent scores that characterize repetition. We also examine diversity in downstream datasets such as instruction tuning. Finally, we show that compression ratio—compressed over original texts size—is a fast, easy to compute score sufficient to capture the information in all token/type ratio related alternatives. But we emphasize text length as an important confounder when assessing diversity: No reliable conclusions can be drawn without taking this into consideration.
@@
Our contributions are as follows. (1) We introduce diversity, a Python package implementing diversity metrics. (2) We host and release source code to a user interface to explore repetition and diversity. (3) We evaluate the convergent validity of existing lexical diversity metrics and highlight compression ratios as efficient measures of diversity.
@@ 2 Related Work
Lack of diversity in text may result from repetition of lengthy strings or owe to subtle distributional patterns (Holtzman et al., 2020; Wiher et al., 2022; meister-etal-2023-locally). We focus on scores that aim to capture overt repetition across outputs, and leave for future work similar analysis of semantic and structural diversity scores (bar-etal-2012-text; shaib-etal-2024-detection). Conditional generation tasks such as image captioning have offered observations regarding the diversity of produced texts. Prior work has shown that models tend to repeat the same text for different contexts in these tasks li-etal-2016-diversity; devlin-etal-2015-language. Self-repetition (Salkar et al., 2022)—exact repetition of the same $n$-gram ($n\geq 4$) across outputs—is a practical way of measuring repetition in lengthy outputs. In such cases repetition is common, especially relative to training data (wang-etal-2023-automated).
We discuss several metrics but it is unclear which of these to use when, and how to efficiently visualize lower diversity in text-only tasks. Further, prior work has shown that human judgments of diversity are difficult to reliably collect. Humans tend to implicitly conflate quality of text with its diversity, and it can be difficult to separate content and lexical diversity in such assessments (tevet-berant-2021-evaluating). We design an interactive tool to allow users to browse highlighted instances of “lower diversity” text (Figure 1 (B)).
@@ 2.1 A Smorgasbord of Text Diversity Scores
Scores used to measure diversity across a corpus of texts derive from two core ideas: Computing average similarity between pairs of outputs produced by the same model for different inputs, and computing variants of token/type ratio. The former are adapted from common approaches to reference-based text generation using standard measures of pairwise similarity; the latter track the diversity of vocabulary measured as the ratio of unique words to total words produced, with outputs from a model concatenated into a single text. We first describe each score, and then present insights regarding their mutual redundancy. All scores are defined for a set of generated texts $D$, each conditioned on its respective input.
Self-BLEU The quality of text in machine translation, summarization, and image captioning is often reported in terms of overlap with a reference text. This idea can be adapted to measure diversity across different outputs by using one generated text as a “reference” and measuring the similarity of other outputs against this. Self-BLEU measures similarity between all text pairs in $D$ using BLEU (Zhu et al., 2018). BLEU could be replaced with other similarity scores, e.g., ROUGE-L or BERTScore. These variants are called homogenization scores and have recently been used to compare the diversity of texts produced under several conditions (Padmakumar and He, 2023).
Homogenization Score (ROUGE-L) All homogenization scores calculate an aggregate similarity across pairs of examples (Equation 1). Here the similarity score of choice is ROUGE-L (lin-och-2004-automatic), which quantifies overlap in terms of longest common sub-sequences between all pairs of text in a corpus instead of the fixed $n$-gram size used in other ROUGE variants:
@@
Homogenization Score (BERTScore) This homogenization score uses BERTScore to measure similarity between documents in Equation 1. Unlike the other scores, it does not count the repetition of specific tokens, but instead uses BERT embeddings to (ideally) capture “semantic” similarity beyond verbatim $n$-gram matches.
Self-repetition Score Self-repetition measures the tendency of LMs to repeat long $n$-grams across different outputs (Salkar et al., 2022).
@@
Where $k$ is the total number of 4-grams in a single document $d$ and $N_{i}$ the number of other summaries in which 4-gram $i$ appears. The final score is the sum of $\text{SRS(}d\text{)}$ divided by $|D|$.
Moving Average Token-Type Ratio The token-type ratio for a text is the unique token count divided by the total token count. This metric captures the repetition of a given word in segments of text and does not explicitly account for longer repeated sequences (Covington and McFall, 2010).
NN-Gram Diversity Score NGD extends the idea of token-type ratio to longer $n$-grams (Padmakumar and He, 2023; meister-etal-2023-locally; li-etal-2023-contrastive), taking a ratio of unique to all $n$-gram counts:
@@ 2.2 Compression Ratios for Diversity
Compression Ratios (CRs) The diversity scores introduced so far are all a function of the number of repeated substrings across outputs. We use gZip to compress the concatenated text of all outputs generated by a model. CR is then the ratio between the size of the compressed file to that of the original. High CRs imply more redundancy:
@@
Part-of-Speech Compression Ratio To capture repeated syntactic patterns, we also compute compression ratios for part-of-speech (POS) tag sequences. We use the NLTK POS tagger  [^5] https://www.nltk.org/api/nltk.tag.html and the Penn Treebank set of 36 tags.
@@ 3.1 Design of the Diversity Package
The diversity package incorporates measures of diversity including lexical, syntactic, and semantic diversity. For lexical/syntactic diversity, we use NLTK (bird-loper-2004-nltk) and SpaCY Honnibal et al. (2020) to tag text with parts of speech and extract $n$-grams. We also include implementations of embedding-based measures (Cox et al., 2021), and QUDSim (Namuduri et al., 2025). Users can install the package via pip (assuming Python 3.10+) and can calculate various diversity metrics over a list of texts as follows:
1from diversity import *
2
3text = ["I enjoy walking with my cute dog...", "I enjoy walking outside with...", "I enjoy jogging on a sunny..."]
4# compression ratios
5cr = compression_ratio(text, ’gzip’)
6cr_pos = compression_ratio(get_pos(text)[1], ’gzip’)
7# homogenization scores
8hs_rougel = homogenization_score(text, ’rougel’)
9hs_bert = homogenization_score(text, ’bertscore’)
10self_bleu = homogenization_score(text, ’bleu’)
11# other
12self_rep = self_repetition_score(text)
13nds = ngram_diversity_score(text, n=4)
14
15# Embedding-based
16rc = remote_clique(text, model="Qwen/Qwen3-Embedding-0.6B", verbose=False)
17
18cd = chamfer_dist(text, model="Qwen/Qwen3-Embedding-0.6B", verbose=False)
19
20# QUDSim
21key = os.environ.get("OPENAI_API_KEY") # requires an OpenAI key
22qud_alignment = qudsim(text, key=key) # list of QUD-based alignments/scores
Users can extract PoS patterns using the extract_patterns function by specifying the $n$-gram length to search for, and the top_n most repeated $n$-grams to return:
1extract_patterns(text, n=5, top_n=100)
This returns a Python dictionary where the keys are the part-of-speech n-grams and the values are the raw text n-grams matching those patterns. The pattern matches are based on the frequency seen across the entire dataset, i.e., a part-of-speech pattern is only a pattern if it appears in more than 2 texts in the original input. Default values consider the top 100 part-of-speech patterns (sorted by frequency).
@@
which returns a list of tuples containing the pattern and matched substring, respectively. Many diversity metrics require pairwise comparisons. With larger datasets, this can become infeasible to compute (see Appendix A). We implement a few methods to increase efficiency: memoization of already computed pairs, and batch pattern searching in the UI.
We also include a function for users to run all metrics and display them in a table to easily compare values:
1compute_all_metrics(corpus=text)
3.2 Metric Visualization and the Web UI
The diversity Web UI offers the same functionality as the package via a no-code UI. Figure 1 shows the main pages of the site: users can begin by (A) either uploading their own text file for analysis or selecting one of the demo datasets provided on the site. Then, the user is prompted to select one of three types of analyses: either (B) to explore part-of-speech patterns, (C) to explore verbatim repeated text, or (D) to measure various diversity metrics of the dataset. Datasets are processed upon upload, and nothing is stored on the backend server aside from the existing demo datasets.
(B) Templates The templates tab allows users to explore extracted part-of-speech $n$-grams in their selected dataset. The left-most column displays pattern length of $n=[2,10]$. The user can then scroll through all of the templates, select some or all, and see the highlighted text in the middle panel corresponding to the template. The templates are assigned a colour when selected to indicate the corresponding matched text. The right-most column provides a reference for all the part-of-speech tags from SpaCY.  [^6] https://spacy.io/usage/linguistic-features The default pattern length is set to $n=4$. Other lengths will load when selected.
(C) Exact Match The exact matches tab allows a user to explore exact text matches in their dataset. The top provides two sliders: the left slider allows the user to set a string length to search for ($n=[2,10]$), and the right the minimum number of documents in which the string must appear ($n=[2,10]$). The minimum document occurrence slider defaults to 2. Once selected, the user can scroll through to see the repeated text in bold, and the full document text in which the string appears, as well as the number of documents.
(D) Diversity Metrics The diversity metrics tabs reports the recommended metrics from our evaluation: Compression Ratio, POS Compression, Self-BLEU, Self-Repetition, and Homogenization with BERTScore. We display these values alongside a guide to the metrics on the right-hand side.